OpenStudy (anonymous):

Traffic signs are regulated by the manual on uniform traffic control devices (MUTCD)the perimeter of a rectangular traffic sign is 100 inches. Also, it's length is 14 inches longer than its width. Find the dimensions of this sign. The width of the sign is___in. The length of the sign is___in.

OpenStudy (anonymous):

you got it ;)

OpenStudy (anonymous):

l=14+w 2(14+w+w)=100 14+ 2w=50

OpenStudy (anonymous):

ok